OpenStudy (anonymous):

Last week, Eva spent seven and three-fifths hours on homework. This week, she spent ten and six-tenths hours. How many more hours did she spend this week? 2 two and three-fifths 3 three and three-fifths

OpenStudy (anonymous):

10 6/10 - 7 3/5 Can you subtract?

OpenStudy (anonymous):

6/ 10 - 3/5 = ( 6 - 6 ) / 10 = 0 10 - 7 = 3 hrs